The MoMi Corporation’s income before interest, depreciation and taxes, was $1.6 million in the year just ended, and it expects that this will grow by 5% per year forever. To make this happen, the firm will have to invest an amount equal to 16% of pretax cash flow each year. The tax rate is 35%. Depreciation was $220,000 in the year just ended and is expected to grow at the same rate as the operating cash flow. The appropriate market capitalization rate for the unleveraged cash flow is 12% per year, and the firm currently has debt of $2 million outstanding. Use the free cash flow approach to calculate the value of the firm and the firm’s equity. (Enter your answer in dollars not in millions.)

Value of the firm $

Value of the firm's equity $
Answer #1

C D E 1 2 Sol) $ $ $ $ $ 1,680,000 231,000 =220000*1.05 1,449,000 =C3-C4 507,150.0 =C5*0.35 941,850.0 =C5-C6 Before-tax cash